OpenStudy (anonymous):

A 50 meter long log is cut into 8 pieces of equal length, with some of the log left over. (The leftover piece is shorter than the 8 equal pieces.) How many meters long is the leftover piece of log?

OpenStudy (callisto):

Number of piece can be cut = 50/8 = 6.25 = 6 (round down) Length of leftover piece = 50 - 8(6) = 50 - 48 = ....

OpenStudy (anonymous):

it's probably over 9000

OpenStudy (valpey):

You may want to make it clear that each of the pieces is an integral number of meters (if that is the case).

OpenStudy (callisto):

Ah... number of piece with equal length can be cut = 50/8 = 6.25 = 6 (round down)

OpenStudy (anonymous):

@petewe you do not know what you are talking about so please do not comment on my stuff

OpenStudy (anonymous):

@Callisto thank you so much.. is the answer 2?

OpenStudy (callisto):

Since number of pieces you have is 7.. 6 with equal length, 1 is shorter.

OpenStudy (callisto):

Yes!!~

OpenStudy (anonymous):

@ callisto so can you explain to me step by step what you did so that I can take notes

OpenStudy (callisto):

Sure :) But I must say that... The answer was correct, though what I've written is not correct. (Forgive me!!)

OpenStudy (anonymous):

you are okay!

OpenStudy (anonymous):

@Callisto are you there?!

OpenStudy (callisto):

A 50 meter long log is cut into 8 pieces of equal length, First, if you divide 50 meter by 8 pieces, you'll know what is the length of each equal part is. 50/8 = 6.25 That means each piece of log cut is 6.25 meters. As @Valpey has mentioned, the assumption made here is that the length of each piece is an integer. So, the length of each piece would reduce to 6 meters. Now, you have 8 pieces of 6 meters each. You can calculate the total length of the equal part. 8 x 6 = 48 So, the leftover part is 50 - 48 = 2m

OpenStudy (callisto):

If the length of each piece is not an integer, you don't need to calculate the leftover.. So, I made that assumption.

OpenStudy (anonymous):

@Callisto thank you so much!

OpenStudy (callisto):

Welcome :) Hope that makes sense to you :)

OpenStudy (anonymous):

yes it makes alot of sense now!